Fermi National Laboratory
 

Accelerator Update for February 7, 2007 - February 9, 2007

 

Wednesday February 7

The day shift began with Operations monitoring the Tevatron (TeV) store 5216 (49.05E30), the Recycler stash (218.8E10), and the Antiproton Source (Pbar) stack (9.3E10), with Booster and Linac in access, with the Main Injector (MI) and Recycler operating normally, with MiniBooNE off, with NuMI off due to the accesses, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

Linac came out of access at 10:16 AM.

The Neutron Therapy Facility (NTF) began using beam for a two-hour study at 10:35 AM.

Operations prepared the Pbar AP1 area for an access at 12:48 PM. They were out of access by 1:30 PM.

Operations began trying to turn on the Booster at 1:31 PM. However, this was slowed by a RF station (BRF9) that still needing its modulator changed and by a problem with the Booster Front End (FE). Operations rebooted the FE at 1:44 PM.

At 2:04 PM, a Tevatron Electron Lens (TEL1) tripped off due to an airflow indication. It reset okay.

At 2:38 PM, Operations reported that a Booster power supply (GMPS) had tripped on an oil level indication and wouldn’t reset. Operators contacted experts.

A Pbar expert began changing out the PLC for the Pbar Stacktail system at 3:01 PM. He was finished in fifteen minutes.

A RF expert reported at 3:11 PM that BRF9 was tied back into the Booster RF system.

Operations reported at 3:52 PM that GMPS1 transformer had been filled with oil.

Operations terminated store 5216 at 4:40 PM.

Meson experts established beam to MTest at 4:52 PM for Switchyard (SY) studies.

Operations had CDF prepared for an access by 5:01 PM.

The SY studies ended at 6:05 PM.

The CDF access ended at 6:53 PM.

Operations put the TeV through a dry squeeze at 7:33 PM.

Operations put the TeV into shot setup at 7:56 PM.

At 8:34 PM, Operations reported that BRF6 had tripped off. It reset okay.

Operations established store 5218 at 10:39 PM with an initial luminosity of 180.42E30.

At 11:53 PM, a Linac Klystron RF station (KRF5) tripped off due to a spark. It reset.

Operations reported at midnight that the Recycler had stashed antiprotons once during the evening shift.

Thursday February 8

The midnight shift began with Operations monitoring store 5218 (136.59E30) and the stash (80.4E10), with Pbar stacking (18.7E10), and with NuMI taking beam.

At 1:11 AM, BRF6 tripped off for the first of six times in five hours. It reset each time.

At 1:19 AM, Operations reported that the Recycler gated cooling turned on again on its own. Operators turned it off.

At 3:24 AM, the TeV store 5218 aborted due to a HFU failure at sector C3. There was no quench. Operations contacted experts.

Operators changed out the HFU at TeV sector C3 at 5:11 AM.

Operations put the TeV through a dry squeeze at 6:03 AM.

Operations put the TeV into shot setup at 6:42 AM.

Operations reported at 8 AM that the Recycler had stashed twice during the midnight shift.

At 8:17 AM, BRF6 tripped off. It reset.

At 9:08 AM, BRF5 tripped off. It reset.

Operations established store 5219 at 9:36 AM with an initial luminosity of 155.18E30.

At 10:02 AM, BRF6 tripped off. It reset.

Operations established beam to SY at 10:16 AM.

NTF completed treatment on its last patient for the day at 12:17 PM.

RF experts began working on BRF6 at 1:20 PM. The work lasted 15 minutes.

At 2:12 PM, BRF6 tripped off. It reset.

Operations reported at 2:47 PM that CUB went into winter mode.

At 3:16 PM, KRF6 caused a beam movement that caused the 400MeV area to trip for the first of two times in three minutes. Linac experts investigated and found a problem with its temperature regulation loop.

At 3:28 PM, BRF6 tripped. It reset.

Operations reported at 4:01 PM that the Recycler had stashed twice during the day shift.

At 5:29 PM, Operations reported that the Recycler electron cooling beam energy was drifting. Operators made adjustments per expert instructions.

At 6:03 PM, Operations reported a problem with the Recycler stochastic cooling system.

SY experts reported at 6:08 PM that they were done with the Meson beam.

At 6:17 PM, BRF6 tripped off. It reset

At 6:31 PM, Operations reported that KRF6 had wandered out of phase.

At 6:33 PM, Operations reported that the Recycler stochastic cooling gating had again turned on. An operator turned it off.

At 6:44 PM, Operations reported that KRF6 was still wandering.

At 6:57 PM, Operations reported that KRF6 was still wandering. Operations contacted an expert. The picture shows the noise that had developed over the last few days on the LCW mixing valve positions.

At 8:05 PM, Operations contacted a Linac expert. The expert reported that the Klystron was losing its phase lock. He suggested that operators reset the water skid and the low-level RF. Operators did this and the station phase-locked.

At 8:34 PM, BRF6 tripped off for the first of many times during the rest of the evening shift. It reset each time.

At 9:58 PM, KRF5 tripped off due to a spark. It reset.

At 11:08 PM, The NuMI beam permit tripped off due to a bad beam position for the first of two times in eight minutes. It reset okay.

Operations reported at 11:59 PM that the Recycler stashed three times during the evening shift.

Friday February 9

The midnight shift began with Operations monitoring store 5219 (37.86E30) and the stash (212.7E10), with Pbar stacking (11.7E10), and with NuMI taking beam.

At 12:21 AM, KRF6 suffered more regulation problems. Resetting didn’t help. Operations continued to investigate.

At 1:16 AM, BRF6 tripped off. It reset.

Operations raised the regulation temperature for KRF6 at 1:23 AM and the station phase-locked.

At 2:01 AM, Operations again raised the regulation temperature for KRF6.

A TeV expert began a TEL2 study at 4:56 AM.

At 5:25 AM, Operations raised the regulation temperature for KRF6.

A TeV expert began a crystal collimator end-of-store study at 6:08 AM.

Operations reported at 8:05 AM that the Recycler stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator wants experts to fix BRF6 and KRF6. He will allow the Recycler, Pbar, and MI experts to conduct studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/12/2007    email Fermilab